Here is another problem I am having trouble with Can someone help? The ratio of SUV's to station wagons sold is 3 to 2. If there were 20 less station wagons sold than SUB's how many of each kind of car were sold?
Thanks for the help.

Answer by checkley71(8403)

SUV/SW
3/2=X/(X-20)
3(X-20)=2X
3X-60=2X
3X-2X=60
X=60 IS THE NUMBER OF SUV'S SOLD.
60-20=40 SW'S SOLD.
PROOF
3/2=60/40
3/2=3/2